A photon has the same momentum as an electron moving with a speed of 1.92×105 m/s. What is the wavelength of the photon?
Answer #1

here ,

speed of photon , v = 1.92 *10^5 m/s

momentum of electron , p = m * v

wavelength of photon = h/mv

wavelength of photon = 6.626 *10^-34/(9.11 *10^-31 * 1.92 *10^5)

wavelength of photon = 3.79 *10^-9 m

the wavelength of the photon is 3.79 *10-9 m